# Mind: Multimodal Shopping Intention Distillation from Large Vision-language Models for E-commerce Purchase Understanding

### v1

@@ Abstract
Improving user experience and providing personalized search results in E-commerce platforms heavily rely on understanding purchase intention. However, existing methods for acquiring large-scale intentions bank on distilling large language models with human annotation for verification. Such an approach tends to generate product-centric intentions, overlook valuable visual information from product images, and incurs high costs for scalability. To address these issues, we introduce Mind, a multimodal framework that allows Large Vision-Language Models (LVLMs) to infer purchase intentions from multimodal product metadata and prioritize human-centric ones. Using Amazon Review data, we apply Mind and create a multimodal intention knowledge base, which contains 1,264,441 intentions derived from 126,142 co-buy shopping records across 107,215 products. Extensive human evaluations demonstrate the high plausibility and typicality of our obtained intentions and validate the effectiveness of our distillation framework and filtering mechanism. Further experiments reveal the positive downstream benefits that Mind brings to intention comprehension tasks and highlight the importance of multimodal generation and role-aware filtering. Additionally, Mind shows robustness to different prompts and superior generation quality compared to previous methods. Our code and data are publicly available at https://github.com/HKUST-KnowComp/MIND_Distillation.
@@
Baixuan Xuthanks:  Equal Contribution, Weiqi Wang, Haochen Shi, Wenxuan Ding, Huihao Jing, Tianqing Fang, Jiaxin Bai, Long Chen, Yangqiu Song Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China bxuan@connect.ust.hk, {wwangbw, longchen, yqsong}@cse.ust.hk
@@ 1 Introduction
Understanding customers’ intentions behind their purchase behaviors remains crucial in E-commerce as it potentially benefits several downstream tasks, such as product recommendation Grbovic et al. (2015); Zhao et al. (2014); Li et al. (2020) and search query answering Zhao et al. (2019); Hirsch et al. (2020). Unlike traditional factual knowledge related to products, intentions are implicit mental states of customers, which typically require commonsense knowledge to understand and reason upon Bratman (1984). For example, in Figure 1, the intentions of purchasing a mouse and a keyboard can be they are very useful to computer users, which is not mentioned either in the customer’s query or products’ metadata. Thus, due to such implicitness, it is infeasible to perform large-scale automatic extraction from text to obtain them.
@@ 2.1 Shopping Intention in E-commerce
Shopping intention is an implicit mental state that motivates purchase-related behaviors from the customer’s perspective Koo and Ju (2010). Various studies have been conducted to examine the impact of consumer shopping intentions on downstream applications Dai et al. (2006); Zhang et al. (2016); Hao et al. (2022). Recently, Ni et al. (2019) suggested using customer reviews to investigate the underlying purchase intentions in consumer purchase behavior and created a large-scale review dataset based on Amazon. Building upon this, Yu et al. (2023) proposed FolkScope, which aims to guide LLMs in generating user co-buy intentions for different product pairs by grounding them in ConceptNet relations Speer et al. (2017). While human evaluations confirmed its effectiveness, Zhou et al. (2024) argued that it not only remains expensive to scale up but also fails to align the resulting shopping intentions with human preferences, which encompass a wide range of factors beyond product properties and similarities. To tackle these issues, in our work, we propose Mind, a framework that undermines online co-buy intentions and aligns better with human perceptions.
@@ 2.2 Multimodal Knowledge Distillation
Since VLMs have yield significant advance recently Liu et al. (2023d); Li et al. (2023); Zhu et al. (2023), distilling domain-specific knowledge from them has become an effective yet cost-friendly trend in multimodal studies Liu et al. (2023c); Lu et al. (2024); Jin et al. (2021). Liu et al. (2023c) proposed a framework that applies self-distillation to stimulate the pre-train process of BERT to improve its performance in E-commerce product understanding tasks. Lu et al. (2024) similarly instructed MiniGPT4 Zhu et al. (2023) to generate user intention form social media posts text and its associated images. Jin et al. (2021) also designed a framework to instruct the student model to imitate teacher model’s behavior, which successfully preserved the teacher model’s capabilities with fewer parameters. In our work, we share the same aspiration and leverage distillation as a tool for data collection that provides downstream benefits in the E-commerce domain. Specifically, we designed a framework to distill E-commerce intentions from LLaVa Liu et al. (2023a) and construct a comprehensive intention knowledge base based on the resulted generations.
@@ 3.2 Source Data Collection
We utilize the Amazon Review Data released by Ni et al. (2019), which contains millions of products from 18 domains. Each product is accompanied by detailed reviews, co-buy records, and metadata, including its product title, features, attributes, and images provided by the retailer. Following Yu et al. (2023), we select products from the Electronics and Clothing, Shoes and Jewelry domains as representative products to demonstrate the effectiveness of Mind. To fit our framework, we filter out products without accessible images that may have been removed from the Amazon platform.
@@ 4.1 Annotation Setup
We hire human annotators from the Amazon Mechanical Turk platform to evaluate the generated intentions. For a generated intention, we task each worker to evaluate four aspects:
@@ 5.1 Evaluation Setup
While existing results show that language models struggle with both tasks, we aim to examine whether Mind can enhance LLMs’ intention understanding capabilities through fine-tuning. Specifically, from all intentions generated by Mind, we transform them into instruction-following format via natural language templates following Zhou et al. (2023). Then, we fine-tune LLAMA2-7B-chat Jiang et al. (2023) and Mistral-7B-Instruct-v0.2 Touvron et al. (2023) on the retrieved data as a type of knowledge injection. They are then evaluated in a zero-shot manner by being prompted to select the most plausible choice for every QA pair in IntentionQA. Accuracy is used as the evaluation metric. More information about the baselines can be found in Appendix D
@@ 5.3.2 Diversity of Mind
Moreover, the semantic diversity of the generated intentions are another flag of the quality as featuring intentions that cover a diverse collection of topics, events, and even mental states makes it more possible to model purchase intentions comprehensively. Thus, following Wang et al. (2024), we sample 30,000 intentions from Mind, extract the nouns in them via dependency parsing, and plot the distribution of the hypernyms of these nouns, matched against Probase Wu et al. (2012), according to their number of occurences. The resulting plot is shown in Figure 3. Remarkably, the intentions generated by Mind display elevated levels of diversity, which signifies the broad semantic coverage of purchasing different products in the generated intentions. We posit that such high semantic diversity may provide implicit benefits to downstream tasks when employing Mind in E-commerce applications.
@@ 5.3.4 Robustness of Mind
According to Chang et al. (2024), generations by LLMs can be significantly impacted by even slight changes in the prompts. This warrants a potential weakness of Mind which heavily relies on prompting in collecting intentions. Hence, we aim to overcome this by proving that intentions generated with modified prompts are generally semantically consistent at high quality. Specifically, we exclude the prompts which explicitly instructing the LVLMs to rely on visual cues from the product images and only retain the prompts that require the LVLMs to generate intentions. Then, 100 product pairs are randomly sampled from Mind to generate intentions utilizing the modified prompts. Finally, the sentence embedding are calculated using SentenceBERT Reimers and Gurevych (2019), and the cosine similarity between each modified intention and its corresponding original intention generated by Mind is derived. The results revealed an average cosine similarity of 0.85 between the intentions generated with modified prompts and those generated by Mind. This high similarity indicates the robustness of intention generation process. Inter-relation intention comparison examples are provide in Appendix E
@@ 5.3.5 Comparisons Against FolkScope
We then compare Mind against FolkScope, the previous state-of-the-art method for large-scale intention acquisition, by analyzing the typicality distribution of intentions across all relations. Specifically, we adopt the same annotation protocols designed by Yu et al. (2023); Wang and Song (2024); Wang et al. (2023b) and transfer our annotation results into a four-point Likert scale Joshi et al. (2015). Then, for each relation, we compute the average typicality scores among all intentions and plot them for comparison, as shown in Figure 5. From the plot, we observe that intentions generated by Mind exhibit higher typicality scores across nearly all relations compared to those generated by FolkScope, which further demonstrates the superiority of Mind. More exemplar-based case comparisons are presented in Appendix F and a relation-wise filter result analysis is attached in Appendix G.
@@ 6 Conclusions
First, Mind generates intention by leveraging several zero-shot prompts without additional exemplars. This decision is made as we observe that few-shot prompts may “guide” LVLM to generate intentions that tend to be similar to the provided exemplars, which harms diversity. However, it remains an open question whether more advanced prompting methods Song et al. (2023); Parnami and Lee (2022) would help in the generation process. It’s also worth noting that the LVLM used in our work may be outdated as new products show up on E-commerce platforms. However, switching LLaVa to more up-to-date LVLMs, preferably pre-trained on E-commerce data, can address this concern. Finally, Mind utilizes an automatically functioning filter as quality control. While we have shown its effectiveness, it remains challenging to effectively regulate the filter mechanism to be either lenient or strict. Further investigation is required to provide insights into the alignment between the values of VLMs and the real world, enhancing our understanding of them.
@@ 7 Acknowledgements
We thank the anonymous reviewers and the area chair for their constructive comments. The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
@@ Appendix C Error Analysis of Filtered Intentions
Appendix D Baseline Backbone
For both tasks, we first incorporate random and majority voting to reflect the characteristics of the benchmark. Five Pre-Trained Language Models (PTLMs) are included: RoBERTa Liu et al. (2019) DeBERTa-v3 He et al. (2023), T0 Sanh et al. (2022), T5 Raffel et al. (2020), and Flan-T5 Chung et al. (2022). Then, performances by five commonsense-injected PTLMs are also reported, including HyKAS Ma et al. (2021), CAR Wang et al. (2023a), VERA Liu et al. (2023b), CANDLE Wang et al. (2024), and VERA-CANDLE. We also report the performances of several LLMs, such as LLaMA2 Touvron et al. (2023), Gemma Mesnard et al. (2024), Mistral Jiang et al. (2023), ChatGPT OpenAI (2022), and GPT-4 OpenAI (2023). For the latter two, we also adopt Chain-of-Thought (CoTWei et al., 2022) and CoT with Self-Consistency (CoT-SCWang et al., 2023c) prompting.
Appendix E Mind Inter-relation Case Study
@@
Item1 Item2 Intentions
They are derived from the same category.
They both are a part of the ’Adult Costume’ category.
They both are a part of the Columbia brand.
They are both used for outdoor activities.
Novelty Video Games Sarcastic Mens Very Funny T Shirt Super Mario All Over Character 100 Score Print Socks They both appeal to gamers and video game enthusiasts.
They both are of high quality and are really cool.
Women’s Exaggerative Vintage Leafs Shape Necklace Sets Plated Oil Drip Rhinestone Flower Necklace Earring Sets They both have a vintage-inspired design and feature colorful flowers and beads
They both have a property of ’High Quality’.
Xbox 360 4gb Kinect Bundle Controller Charger - Xbox 360 They both cater to the needs of Xbox 360 gamers.
They both are made of plastic.
Table 5: Case studies of purchase intentions generated by Mind and FolkScope. Intentions generated by Mind are highlighted in blue and those generated by FolkScope are marked in green.
Appendix F Mind Against FolkScope Case Study
Aside from empirical analyses, we also show the advantages of Mind over FolkScope through additional case studies to highlight key benefits of Mind. To this end, we randomly selected 7 pairs of co-buy products and compared the intentions generated by both frameworks, as shown in Table 5. Our findings from the table indicate that Mind-generated intentions exhibit a stronger focus on the usage and functionalities that potentially fulfill customers’ needs and intentions when purchasing these products. Conversely, intentions generated by FolkScope tend to be biased towards properties and features that can be easily inferred from the product titles, which are of lesser interest to customers’ shopping intentions. Take the second row in Table 5 as an example. The intentions both are “Women’s Shoes” and “Women’s Handbags” generated by FolkScope merely represent an aggregation of the product categories for the two items. In contrast, Mind produces intentions such as looking for stylish and functional combination for daily activities , which better captures a customer’s intention when shopping for both products. This example further reinforces our previous conclusions that Mind can generate intentions that are more human-centric and better reflect the customers’ intentions as mental states.

### v3

@@ Abstract
Improving user experience and providing personalized search results in E-commerce services heavily rely on understanding purchase intention. However, existing methods for acquiring large-scale intentions bank on distilling large language models with human annotation for verification. Such an approach tends to generate product-centric intentions, overlook valuable visual information from product images, and incurs high costs for scalability. To address these issues, we introduce Mind, a multimodal framework that allows Large Vision-Language Models (LVLMs) to infer purchase intentions from multimodal product metadata and prioritize human-centric ones. Using Amazon Review data, we apply Mind and create a multimodal intention knowledge base, which contains 1,264,441 intentions derived from 126,142 co-buy shopping records across 107,215 products. Extensive human evaluations demonstrate the high plausibility and typicality of our obtained intentions and validate the effectiveness of our distillation framework and filtering mechanism. Further experiments reveal the positive downstream benefits that Mind brings to intention comprehension tasks and highlight the importance of multimodal generation and role-aware filtering. Additionally, Mind shows robustness to different prompts and superior generation quality compared to previous methods. Our code and data are publicly available at https://github.com/HKUST-KnowComp/MIND_Distillation.
@@
Baixuan Xu1thanks:  Equal Contribution , Weiqi Wang1, Haochen Shi1, Wenxuan Ding1, Huihao Jing1, Tianqing Fang1, Jiaxin Bai1, Xin Liu2, Changlong Yu2, Zheng Li2, Chen Luo2, Qingyu Yin2, Bing Yin2, Long Chen1, Yangqiu Song1 1Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China 2Amazon.com Inc, Palo Alto, CA, USA bxuan@connect.ust.hk, {wwangbw, longchen, yqsong}@cse.ust.hk {xliucr, changlyu, amzzhe, cheluo, qingyy, alexbyin}@amazon.com
@@ 1 Introduction
Understanding customers’ intentions behind their purchase behaviors remains crucial in E-commerce as it potentially benefits several downstream tasks, such as product recommendation Grbovic et al. (2015); Zhao et al. (2014); Li et al. (2020) and search query answering Zhao et al. (2019); Hirsch et al. (2020); Bai et al. (2023). Unlike traditional factual knowledge related to products, intentions are implicit mental states of customers, which typically require commonsense knowledge to understand and reason upon Bratman (1984). For example, in Figure 1, the intentions of purchasing a mouse and a keyboard can be they are very useful to computer users, which is not mentioned either in the customer’s query or products’ metadata. Thus, due to such implicitness, it is infeasible to perform large-scale automatic extraction from text to obtain them.
@@ 2.1 Shopping Intention in E-commerce
Shopping intention is an implicit mental state that motivates purchase-related behaviors from the customer’s perspective Koo and Ju (2010). Various studies have been conducted to examine the impact of consumer shopping intentions on downstream applications Dai et al. (2006); Zhang et al. (2016); Hao et al. (2022); Wang et al. (2024a). Recently, Ni et al. (2019) suggested using customer reviews to investigate the underlying purchase intentions in consumer purchase behavior and created a large-scale review dataset based on Amazon. Building upon this, Yu et al. (2023) proposed FolkScope, which aims to guide LLMs in generating user co-buy intentions for different product pairs by grounding them in ConceptNet relations Speer et al. (2017). While human evaluations confirmed its effectiveness, Zhou et al. (2024) argued that it not only remains expensive to scale up but also fails to align the resulting shopping intentions with human preferences, which encompass a wide range of factors beyond product properties and similarities. To tackle these issues, in our work, we propose Mind, a framework that undermines online co-buy intentions and aligns better with human perceptions.
@@ 2.2 Multimodal Knowledge Distillation
Since VLMs have yield significant advance recently Liu et al. (2023d); Li et al. (2023); Zhu et al. (2023); Chan et al. (2024); Zong et al. (2023), distilling domain-specific knowledge from them has become an effective yet cost-friendly trend in multimodal studies Liu et al. (2023c); Lu et al. (2024); Jin et al. (2021). Liu et al. (2023c) proposed a framework that applies self-distillation to stimulate the pre-train process of BERT to improve its performance in E-commerce product understanding tasks. Lu et al. (2024) similarly instructed MiniGPT4 Zhu et al. (2023) to generate user intention form social media posts text and its associated images. Jin et al. (2021) also designed a framework to instruct the student model to imitate teacher model’s behavior, which successfully preserved the teacher model’s capabilities with fewer parameters. In our work, we share the same aspiration and leverage distillation as a tool for data collection that provides downstream benefits in the E-commerce domain. Specifically, we designed a framework to distill E-commerce intentions from LLaVa Liu et al. (2023a) and construct a comprehensive intention knowledge base based on the resulted generations.
@@ 3.2 Source Data Collection
We utilize the Amazon Review Data released by Ni et al. (2019), which contains millions of products from 18 domains. Each product is accompanied by detailed reviews, co-buy records, and metadata, including its product title, features, attributes, and images provided by the retailer. Following Yu et al. (2023), we select products from the Electronics and Clothing, Shoes and Jewelry domains as representative products to demonstrate the effectiveness of Mind. It is important to note that while this work focuses on only two domains, the Mind framework itself is not limited to these. It can be applied to any product domain, allowing for seamless deployment to undermine co-buy intentions across a wide range of products beyond those selected. To fit our framework, we filter out products without accessible images that may have been removed from the Amazon service.
@@ 4.1 Annotation Setup
We hire human annotators from the Amazon Mechanical Turk service to evaluate the generated intentions. For a generated intention, we task each worker to evaluate four aspects:
@@ 5.1 Evaluation Setup
While existing results show that language models struggle with both tasks, we aim to examine whether Mind can enhance LLMs’ intention understanding capabilities through fine-tuning. Specifically, from all intentions generated by Mind, we transform them into instruction-following format via natural language templates following Zhou et al. (2023). Then, we fine-tune LLAMA2-7B-chat Jiang et al. (2023) and Mistral-7B-Instruct-v0.2 Touvron et al. (2023) on the retrieved data as a type of knowledge injection. Specifically, we adopt LLaMA-Factory Zheng et al. (2024) through our fine-tune process, maintaining default hyperparameters and a LoRA rank of 64. The fine-tune instruction is attached in Appendix D. All experiments are performed on a Linux machine with eight NVIDIA V100 GPUs. They are then evaluated in a zero-shot manner by being prompted to select the most plausible choice for every QA pair in IntentionQA. The zero-shot evaluation setup could be found at Appendix E. Accuracy is used as the evaluation metric.
5.2 Baseline Backbone
For both tasks, we first incorporate random and majority voting to reflect the characteristics of the benchmark. Five Pre-Trained Language Models (PTLMs) are included: RoBERTa Liu et al. (2019) DeBERTa-v3 He et al. (2023), T0 Sanh et al. (2022), T5 Raffel et al. (2020), and Flan-T5 Chung et al. (2022). Then, performances by five commonsense-injected PTLMs are also reported, including HyKAS Ma et al. (2021), CAR Wang et al. (2023a), VERA Liu et al. (2023b), CANDLE Wang et al. (2024b), and VERA-CANDLE. We also report the performances of several LLMs, such as LLaMA2 Touvron et al. (2023), Gemma Mesnard et al. (2024), Mistral Jiang et al. (2023), ChatGPT OpenAI (2022), and GPT-4 OpenAI (2023). For the latter two, we also adopt Chain-of-Thought (CoTWei et al., 2022) and CoT with Self-Consistency (CoT-SCWang et al., 2023c) prompting.
@@ 5.4.2 Diversity of Mind
Moreover, the semantic diversity of the generated intentions are another flag of the quality as featuring intentions that cover a diverse collection of topics, events, and even mental states makes it more possible to model purchase intentions comprehensively. Thus, following Wang et al. (2024b), we sample 30,000 intentions from Mind, extract the nouns in them via dependency parsing, and plot the distribution of the hypernyms of these nouns, matched against Probase Wu et al. (2012), according to their number of occurences. The resulting plot is shown in Figure 3. Remarkably, the intentions generated by Mind display elevated levels of diversity, which signifies the broad semantic coverage of purchasing different products in the generated intentions. We posit that such high semantic diversity may provide implicit benefits to downstream tasks when employing Mind in E-commerce applications.
@@ 5.4.4 Robustness of Mind
According to Chang et al. (2024), generations by LLMs can be significantly impacted by even slight changes in the prompts. This warrants a potential weakness of Mind which heavily relies on prompting in collecting intentions. Hence, we aim to overcome this by proving that intentions generated with modified prompts are generally semantically consistent at high quality. Specifically, we exclude the prompts which explicitly instructing the LVLMs to rely on visual cues from the product images and only retain the prompts that require the LVLMs to generate intentions. Then, 100 product pairs are randomly sampled from Mind to generate intentions utilizing the modified prompts. Finally, the sentence embedding are calculated using SentenceBERT Reimers and Gurevych (2019), and the cosine similarity between each modified intention and its corresponding original intention generated by Mind is derived. The results revealed an average cosine similarity of 0.85 between the intentions generated with modified prompts and those generated by Mind. This high similarity indicates the robustness of intention generation process. Inter-relation intention comparison examples are provide in Appendix F
@@ 5.4.5 Comparisons Against FolkScope
We then compare Mind against FolkScope, the previous state-of-the-art method for large-scale intention acquisition, by analyzing the typicality distribution of intentions across all relations. Specifically, we adopt the same annotation protocols designed by Yu et al. (2023); Wang and Song (2024); Wang et al. (2023b, 2024c) and transfer our annotation results into a four-point Likert scale Joshi et al. (2015). Then, for each relation, we compute the average typicality scores among all intentions and plot them for comparison, as shown in Figure 5. From the plot, we observe that intentions generated by Mind exhibit higher typicality scores across nearly all relations compared to those generated by FolkScope, which further demonstrates the superiority of Mind. More relation-wise filter result analysis is attached in Appendix G.
Item1 Item2 Intentions
They are derived from the same category.
They both are a part of the ’Adult Costume’ category.
They both are a part of the Columbia brand.
They are both used for outdoor activities.
Table 4: Case studies of purchase intentions generated by Mind and FolkScope. Intentions generated by Mind are highlighted in purple and those generated by FolkScope are marked in orange.
5.4.6 Mind Against FolkScope Case Study
Aside from empirical analyses, we also show the advantages of Mind over FolkScope through additional case studies to highlight key benefits of Mind. To this end, we randomly selected 5 pairs of co-buy products and compared the intentions generated by both frameworks, as shown in Table 4. Our findings from the table indicate that Mind-generated intentions exhibit a stronger focus on the usage and functionalities that potentially fulfill customers’ needs and intentions when purchasing these products. Conversely, intentions generated by FolkScope tend to be biased towards properties and features that can be easily inferred from the product titles, which are of lesser interest to customers’ shopping intentions. Take the second row in Table 4 as an example. The intentions both are “Women’s Shoes” and “Women’s Handbags” generated by FolkScope merely represent an aggregation of the product categories for the two items. In contrast, Mind produces intentions such as looking for stylish and functional combination for daily activities , which better captures a customer’s intention when shopping for both products. This example further reinforces our previous conclusions that Mind can generate intentions that are more human-centric and better reflect the customers’ intentions as mental states.
@@ 6 Conclusions
First, Mind generates intention by leveraging several zero-shot prompts without additional exemplars. This decision is made as we observe that few-shot prompts may “guide” LVLM to generate intentions that tend to be similar to the provided exemplars, which harms diversity. However, it remains an open question whether more advanced prompting methods Song et al. (2023); Parnami and Lee (2022) would help in the generation process. It’s also worth noting that the LVLM used in our work may be outdated as new products show up on E-commerce stores. However, switching LLaVa to more up-to-date LVLMs, preferably pre-trained on E-commerce data, can address this concern. Finally, Mind utilizes an automatically functioning filter as quality control. While we have shown its effectiveness, it remains challenging to effectively regulate the filter mechanism to be either lenient or strict. Further investigation is required to provide insights into the alignment between the values of VLMs and the real world, enhancing our understanding of them.
@@ 7 Acknowledgements
We thank the anonymous reviewers and the area chair for their constructive comments. The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from Amazon.
@@ Appendix C Error Analysis of Filtered Intentions
Appendix D Fine-tuning Instructions
The instruction adopted during the fine-tuning process is attached below:
Prompt Template for Fine-tuning Question: Q: customer buys <product 1> and <product 2>. What is the most likely intention for buying them? Answer:<intention>.
Where <product 1> and <product 2> are the products co-bought together and the <intention> refers to the co-buy intention retreived from Mind.
Appendix E Zero-shot Evaluation Setup
For the zero-shot evaluation process, we adopt the same hyperparameter setting as the default model provided bu HuggingFace. The Zero-shot evaluation prompt are as below:
Evaluation Prompt for Intention Understanding Question: A customer buys <product 1> and <product 2>. What is the most likely intention for buying them? Candidate Answers: A. because <intention 1>. B. because <intention 2>. C. because <intention 3>. D. because <intention 4>. Answer A or B or C or D only without any other word.
The <product 1> and <product 2> are products co-bought together and the 4 intentions in the candidate answers are four intentions originate from FolkScope. According to the restriction of most likely, only one intention could be the correct answer.
Evaluation Prompt for Intention Utilization Question: A customer buys <product>, because <intention>. What is the customer’s most probable additional purchase? Candidate Answers: A. <product 1> B. <product 2> C. <product 3> D. <product 4> Answer A or B or C or D only without any other word.
The intention refers to the co-buy intention retreived from FolkScope. The 4 products are the products recorded in FolkScope. According to the restriction of most probable, only one product could be the correct answer.
Appendix F Mind Inter-relation Case Study